OpenStudy (anonymous):

Rewrite the exponential equation in logarithmic form: 3^b=17

OpenStudy (anonymous):

Theory : log(a^b) = b*log(a) Using the above theory ; 3^b=17 => log(3^b) = log(17) => b*log(3) = log(17) => b = log(17)/log(3)